# envname-P envname#1

### v2

@@ Appendix D Cluster-aware Spatial Optimization Supplementary
We present the implementation details of algorithms involved in cluster-aware spatial optimization.
D.1. SolveTSP
‘SolveTSP’ implements simulated annealing algorithm for efficiently solving the TSP problem with a large set of candidates. Simulated annealing is a classic metaheuristic approach where the model iteratively proposes new solution and replaces the current solution if a certain condition is satisfied until the temperature goes down to zero. We detail the implementation for simulated annealing in Algo 3.
@@ D.1. SolveTSP
1:procedure SimulatedAnnealing(cities, $\small T_{init}$, $\small T_{min}$, $\small\alpha$)
2:  $\small solution\leftarrow\text{RandomSolution}(cities)$
3:  $\small T\leftarrow T_{init}$
4:  while $\small T>T_{min}$ do
5:   $\small newSolution\leftarrow\text{Neighbor}(solution)$
6:   $\small costDifference\leftarrow\text{Cost}(newSolution)-\text{Cost}(solution)$
7:   if $\small costDifference<0\text{ or }\exp(-costDifference/T)>\text{Random}()$ then
8:     $\small solution\leftarrow newSolution$
@@
10:   $\small T\leftarrow\alpha\times T$
@@
  • $\small T_{init}$ and $\small T_{min}$: The initial and minimum temperatures for the SA algorithm. In our implementation, $\small T_{init}$ is set to 5000 and $\small{T_{min}}$ is set to 0.
  • $\small\alpha$: The cooling rate that determines how fast the temperature decreases. In our implementation, $\small\alpha$ is set to 0.99.
D.2. SolveTSPWithEndpoints
@@
2:Solve the following linear program:
Minimize: $\small\displaystyle\min\sum_{i\neq j}x_{ij}\cdot dist[i][j]$
[…]
Subject to: $\small\displaystyle\sum_{i\neq k}x_{ik}=1,\quad\forall k\neq s,e$
$\small\displaystyle\sum_{i\neq k}x_{ki}=1,\quad\forall k\neq s,e$
$\small\displaystyle{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Add constraints for source node and sink node}}$
$\small\displaystyle\sum_{i\neq s}x_{si}=1$
$\small\displaystyle\sum_{i\neq s}x_{is}=0$
$\small\displaystyle\sum_{i\neq e}x_{ie}=1$
$\small\displaystyle\sum_{i\neq e}x_{ei}=0$
$\small\displaystyle{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Eliminates all subtours}}$
$\small\displaystyle\sum_{i\in S}\sum_{j\notin S,j\neq i}x_{ij}\leq|S|-1,\quad\forall S\subset\{1,\ldots,n\},S\neq\emptyset,S\neq\{1,\ldots,n\}$
$\small\displaystyle{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Add binary variable constraints}}$
$\small\displaystyle x_{ij}\in\{0,1\},\quad\forall i,j$
@@
In each cluster, the dataset typically comprises a limited set of candidate points. Consequently, the prioritization shifts towards optimizing the accuracy of the resultant solution, rather than focusing solely on the computational efficiency. To address the Traveling Salesman Problem (TSP) with predetermined starting and ending points, we adopt a linear programming (LP) methodology. We detail the formulation of linear program in Alg. 4.
@@ Appendix E Overview of the POI extraction pipeline
As illustrated in Sec. 4.2, we design a pipeline to automatically extract POIs and relevant information from user-generated content on various social media platforms. The pipeline consists of the following steps:
@@
  • Use Automatic Speech Recognition (ASR) to obtain transcription from the video, and Optical Character Recognition (OCR) to extract text from the images. Merge them with the original text to obtain the post information.
@@
  • Use map service API to look up the extracted POI names and obtain the coordinates and extract POI names, similar to the evaluation pipeline in Sec.5.1.2.
@@
The prompt for extracting POI names and locations is provided below.
Prompt for Extracting POI Names and Locations # Guidelines ## Task Background Your task is to identify and extract mentioned cations in the posts/travelogues provided by users to help them quickly find these places on a map. Now, based on the content of the post and in context, carry out the extraction and description of Points of Interest (POIs) mentioned in the post. Focus primarily on places that can be visited, rather than merely on place names. ## Notes on Handling POIs 1. Comprehensive Definition of POI: Typically used to describe a specific geographical location or site, such as restaurants, hotels, streets, attractions, museums, bars, cafes, malls, etc. These locations or sites may have specific value or interest to users or travelers. 2. Characteristics of POI: Specific places recommended or mentioned in the post that are usable for dining, entertainment, etc. 3. Specificity: A POI refers to a specific, particular place, not a broad geographical area or city name. 4. Uniqueness: When a text is separated by symbols like ”/”, ”&”, ”,”, for example, Julu Road/Tianzifang”, it often represents two POIs, in this case, Julu Road and Tianzifang should be extracted separately. 5. Examples of POI: Specific restaurants, performance venues, attractions, shops, streets, etc. 6. Non-POI Examples: Collections of places, food names, types of cuisine, performance groups, exhibition events, etc. ## Post Structure Title: The posts title. Text: The main body content of the post. Text in the images: text recognized from the images. Transcribed text: text transcribed from the video. ## Task Process 1. Extraction: Based on your reasoning, judgment, and knowledge, extract all mentioned POIs from the post. 2. Verification: In the context of each POI, ensure all POIs fit the definition and are specific places. 3. Address Information: In the context of each POI, find related address information that can be searched on a map, such as ”158 Julu Road, Shanghai.” 4. Handling No Information: If no location information is available, return an empty POI list: {{}}. 5. Formatting: Organize information into the specified JSON structure. ## Output Format ### Specific Format {{ POI Name”: Related Address Information for the POI }} ### Examples Example 1: If the original post mentions Lao Nong Tang Noodle Shop in Luxi: A time-honored noodle shop that appears on Shanghais must-eat list all year round!”, the output for this POI should be {{ Lao Nong Tang Noodle Shop in Luxi”: null }} Example 2: If the original post mentions Red Baron (Jianye District Wentiyi Road branch) Looking around, the most striking red on the entire Wentiyi Road, seamlessly blending with Mixue Bingcheng”, the output for this POI should be {{ Red Baron (Jianye District Wentiyi Road branch)”: null }} ## Output Standards - The output is a dictionary, with keys being the POI names and values being the related address information for the POI. If address information is missing, please use null to fill in. - Ensure the output is in valid JSON format and can be parsed by Python json.loads. ## Task Start Please begin processing the post content: “‘{post_info}“‘. Note: Ensure the return format follows {{Point of Interest Name: Related Address Information}}. Ensure it can be json.loads parsed.
@@
Prompt for Generating POI Descriptions {post_info} Based on the content of the above post, please write out the reasons for recommending each location to tourists in the following list. Consider what can be done at this location, its features, and why it is fun. If the original post lacks information, you may appropriately supplement based on your knowledge, but please ensure brevity. The related information for each point should not exceed 30 words. The results should be output in JSON format, specifically in the form {{Place Name: Information related to the place from the original post}}, where Information related to the place from the original post should be a sentence or phrase, like a string. If a place does not have any relevant information, fill in the description related to the place from the original post with null”.
We execute an automated process to extract POIs from the most recent trending posts and update a comprehensive POl database. At a regular interval of 24 hours, we obtain recent trending travel-related posts across multiple cities on social media platforms, and run the above pipeline to extract POI names, locations and descriptions to maintain the database.
@@ Appendix F Prompts
In this section, we provide all the prompts for LLMs used in our method and evaluation.
F.1. Prompt for Decomposing User Requests
@@
Prompt for Decomposing user requests Please help me break down a user request into multiple independent requirements, each including both positive and negative requirements. Return the results in the following format directly based on the **User Request** given, without writing any code. ### Output Format: Return a list, where each item is a dictionary representing an independent requirement, with the following key-value pairs: - **pos**: The positive requirement, representing what the user wants, excluding any negative requirements. - **neg**: The negative requirement, generally what the user does not want, dislikes, or refuses. All negative requirements must be captured in this field, for example, non-spicy should extract spicy”, dont want crowded places should extract crowded”, hate noisy should extract noisy”. - **mustsee**: Indicates whether this requirement represents a specific place name. If so, this field is true‘, otherwise, it is false‘. - **type**: Indicates whether the requirement is for a place or an itinerary”, with place having sub-types place”, starting point”, and ending point”. Overall, this field can have the values place”, starting point”, ending point”, or itinerary”. - Your return should be a list in the following format: [ {{ pos”: positive requirement”, (excluding negative requirements) neg”: negative requirement (whats not wanted, disliked, refused, not wanting to go or see, any negated requirement), mustsee”: true (whether its a must-see point, all specific places should be set to true), type”: place }}, ] - The **positive requirement** must not be empty, and it must not include any negative requirements. All negative requirements should be summarized in the value of the neg field. - Set to null in cases where there are no **negative requirements** for a specific place. - Sometimes users only describe what they do not want (negative requirements), in such cases, you should summarize a **positive requirement** based on the **negative requirement**. For example, if a user says dont want spicy food’, the output should include: pos corresponding to food”, neg corresponding to spicy”. - Independent requirements must have specific descriptions or demands to be considered a requirement, for example, recommend a route does not count as an independent requirement. - mustsee must be a specific place name, not a general term. - If a place is definitively a starting point or ending point”, then the value of the type field should be starting point or ending point”, respectively. Starting points and ending points are must-see points, with the mustsee field set to true. - A place can only be considered as a starting point or ending point if it is a specific attraction or location, and there can only be at most one starting point and one ending point”. - The return should not include any other content. ### Example Outputs: Example 1: User Request: I want to start by visiting Sinan Mansions, then find something fun to do nearby, and I dont want it to be crowded Output: [ {{ pos”: Sinan Mansions”, neg”: null, mustsee”: true, type”: starting point }}, {{ pos”: fun places near Sinan Mansions”, neg”: crowded”, mustsee”: false, type”: place }} ] ### mustsee Field Assignment Examples mustsee true for specific place names: Hualian Supermarket”, Old Mac Coffee Shop”, Wukang Mansion”, Nanluoguxiang”, mustsee false for general place names: supermarket”, milk tea shop”, bar”, coffee”, ### Output Guidelines - Return a list, each item in the list is a dictionary containing pos”, neg”, mustsee”, and type key-value pairs. - Return as a JSON List. - The list can be empty; if empty, just return a JSON list. - The output should not include any other information, ensuring it can be parsed by json.loads. ### User Request {user_req} ### Task Overview Your task is to analyze and break down the **User Request** into independent requirements and return them. 1. First, separate the different independent requirements, breaking down each into positive and negative requirements. 2. Positive requirements should only include what the user wants, and negative requirements should only include what the user does not want. 3. For each independent requirement, refer to **mustsee field assignment examples** to assign a value to the mustsee field, analyzing whether the **positive requirement** is a specific place name. If so, set mustsee to true, otherwise set it to false. 4. Refer to the **examples** and **output format** to complete the other fields. #### Notes: - Do not include duplicate independent requirements; ensure each independent requirement corresponds to different key points in the users needs. - Itinerary requirements should be for the whole itinerary, such as including several places, approximate time, etc., all others are place requirements. - The type field can only be one of [”place”, itinerary”, starting point”, ending point”]. - All attractions must be completely separated, such as Nanluoguxiang and Drum Tower must be split into Nanluoguxiang and Drum Tower as two requirements. Now, based on the **user Request**, refer to the **example outputs**, and return according to the **output guidelines** and **output format**.
F.2. Prompt for Indicating Travel Time of an Itinerary
@@
Prompt for Travel Time Indication Please play the role of a top AI Travel Time Planning Assistant. Your job is to determine the time needed for a days itinerary based on the user request. If the user request is empty, please default to [”4”]. ## Task Overview Your task is to return the required time for a days itinerary based on given the user request. If the user request mention specific time constraints for the route, return the itinerary time directly as per the users request, up to a maximum of 8 hours (return [”8”] if it exceeds 8 hours). Please return the itinerary time directly based on the user request, no need to write any code. ## User Request {user_reqs} ## Input-Output Examples - **Example 1**: - **User Request**: Id like to visit a museum, enjoy authentic cuisine, and experience nightlife.” - **Output**: [”8”] - **Example 2**: - **User Request**: I want to tour historical buildings and take in the city views - **Output**: [”6”] - **Example 3** (In this example, the user specifies approximately **five hours** for the route): - **User Request**: I plan to explore the Huangpu River and Yu Garden for about five hours.” - **Output**: [”5”] ## Output Specifications - Return a list of length 1, containing a single integer representing the required itinerary duration (in hours). The value range is 1 to 8. - Return as a JSON List with only one element inside. - The list can be empty; please only return a JSON list. - Ensure your output contains no additional information and can be parsed by json.loads. Now, based on the **User Request**, return the time required for a days itinerary according to the **Output Specifications**.
@@
F.3. Prompt for Identifying the Start POI

### v4

@@ Appendix D Cluster-aware Spatial Optimization Supplementary
We present the details of the implementation of algorithms involved in cluster-aware spatial optimization.
D.1 SolveTSP
@@
1:procedure SimulatedAnnealing(cities, $T_{init}$, $T_{min}$, $\alpha$)
2:  $solution\leftarrow\text{RandomSolution}(cities)$
3:  $T\leftarrow T_{init}$
4:  while $T>T_{min}$ do
5:   $newSolution\leftarrow\text{Neighbor}(solution)$
6:   $costDifference\leftarrow\text{Cost}(newSolution)-\text{Cost}(solution)$
7:   if $costDifference<0\text{ or }\exp(-costDifference/T)>\text{Random}()$ then
8:     $solution\leftarrow newSolution$
@@
10:   $T\leftarrow\alpha\times T$
@@
‘SolveTSP’ implements a simulated annealing algorithm for efficiently solving the TSP problem with a large set of candidates. Simulated annealing is a classic metaheuristic approach where the model iteratively proposes a new solution and replaces the current solution if a certain condition is satisfied until the temperature goes down to zero. We detail the implementation for simulated annealing in Algo 3.
@@
  • $T_{init}$ and $T_{min}$: The initial and minimum temperatures for the SA algorithm. In our implementation, $T_{init}$ is set to 5000 and $T_{min}$} is set to 0.
  • $\alpha$: The cooling rate that determines how fast the temperature decreases. In our implementation, $\alpha$ is set to 0.99.
D.2 SolveTSPWithEndpoints
In each cluster, the dataset typically comprises a limited set of candidate points. Consequently, the prioritization shifts towards optimizing the accuracy of the resultant solution rather than focusing solely on computational efficiency. To address the Traveling Salesman Problem (TSP) with predetermined starting and ending points, we adopt a linear programming (LP) methodology. We detail the formulation of the linear program in Alg. 4.
@@ D.2 SolveTSPWithEndpoints
1:distdist, start_pointstart\_point, end_pointend\_point
2:Solve the following linear program:
Minimize: $\min\sum_{i\neq j}x_{ij}\cdot dist[i][j]$
@@
Subject to: $\sum_{i\neq k}x_{ik}=1,\quad\forall k\neq s,e$
$\sum_{i\neq k}x_{ki}=1,\quad\forall k\neq s,e$
${\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Add constraints for source node and sink node}}$
$\sum_{i\neq s}x_{si}=1$
$\sum_{i\neq s}x_{is}=0$
$\sum_{i\neq e}x_{ie}=1$
$\sum_{i\neq e}x_{ei}=0$
${\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Eliminates all subtours}}$
$\sum_{i\in S}\sum_{j\notin S,j\neq i}x_{ij}\leq|S|-1,\quad\forall S\subset\{1,\ldots,n\},S\neq\emptyset,S\neq\{1,\ldots,n\}$
${\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Add binary variable constraints}}$
$x_{ij}\in\{0,1\},\quad\forall i,j$
@@ Appendix E Overview of the POI extraction pipeline
Prompt for Extracting POI Names and Locations # Guidelines ## Task Background Your task is to identify and extract mentioned cations in the posts/travelogues provided by users to help them quickly find these places on a map. Now, based on the content of the post and in context, carry out the extraction and description of Points of Interest (POIs) mentioned in the post. Focus primarily on places that can be visited, rather than merely on place names. ## Notes on Handling POIs 1. Comprehensive Definition of POI: Typically used to describe a specific geographical location or site, such as restaurants, hotels, streets, attractions, museums, bars, cafes, malls, etc. These locations or sites may have specific value or interest to users or travelers. 2. Characteristics of POI: Specific places recommended or mentioned in the post that are usable for dining, entertainment, etc. 3. Specificity: A POI refers to a specific, particular place, not a broad geographical area or city name. 4. Uniqueness: When a text is separated by symbols like "/", "&", ",", for example, "Julu Road/Tianzifang", it often represents two POIs, in this case, "Julu Road" and "Tianzifang" should be extracted separately. 5. Examples of POI: Specific restaurants, performance venues, attractions, shops, streets, etc. 6. Non-POI Examples: Collections of places, food names, types of cuisine, performance groups, exhibition events, etc. ## Post Structure Title: The posts title. Text: The main body content of the post. Text in the images: text recognized from the images. Transcribed text: text transcribed from the video. ## Task Process 1. Extraction: Based on your reasoning, judgment, and knowledge, extract all mentioned POIs from the post. 2. Verification: In the context of each POI, ensure all POIs fit the definition and are specific places. 3. Address Information: In the context of each POI, find related address information that can be searched on a map, such as "158 Julu Road, Shanghai." 4. Handling No Information: If no location information is available, return an empty POI list: {{}}. 5. Formatting: Organize information into the specified JSON structure. ## Output Format ### Specific Format {{ "POI Name": "Related Address Information for the POI" }} ### Examples Example 1: If the original post mentions "Lao Nong Tang Noodle Shop in Luxi: A time-honored noodle shop that appears on Shanghais must-eat list all year round!", the output for this POI should be {{ "Lao Nong Tang Noodle Shop in Luxi": null }} Example 2: If the original post mentions "Red Baron (Jianye District Wentiyi Road branch) Looking around, the most striking red on the entire Wentiyi Road, seamlessly blending with Mixue Bingcheng", the output for this POI should be {{ "Red Baron (Jianye District Wentiyi Road branch)": null }} ## Output Standards - The output is a dictionary, with keys being the POI names and values being the related address information for the POI. If address information is missing, please use "null" to fill in. - Ensure the output is in valid JSON format and can be parsed by Python json.loads. ## Task Start Please begin processing the post content: “‘{post_info}“‘. Note: Ensure the return format follows {{Point of Interest Name: Related Address Information}}. Ensure it can be json.loads parsed.
The prompt for extracting POI names and locations is provided above. As illustrated in Sec. 3.2, we design a pipeline to automatically extract POIs and relevant information from user-generated content on various social media platforms. The pipeline consists of the following steps:
@@
  • Use automatic speech recognition to obtain transcription from the video and optical character recognition (OCR) to extract text from the images. Merge them with the original text to obtain the post information.
@@
  • Use map service API to look up the extracted POI names, obtain the coordinates, and extract POI names, similar to the evaluation pipeline in Sec.C.
@@
Prompt for Generating POI Descriptions {post_info} Based on the content of the above post, please write out the reasons for recommending each location to tourists in the following list. Consider what can be done at this location, its features, and why it is fun. If the original post lacks information, you may appropriately supplement based on your knowledge, but please ensure brevity. The related information for each point should not exceed 30 words. The results should be output in JSON format, specifically in the form {{Place Name: Information related to the place from the original post}}, where "Information related to the place from the original post" should be a sentence or phrase, like a string. If a place does not have any relevant information, fill in the description related to the place from the original post with "null".
We execute an automated process to extract POIs from the most recent trending posts and update a comprehensive POl database. At a regular interval of 24 hours, we obtain recent trending travel-related posts across multiple cities on social media platforms and run the above pipeline to extract POI names, locations, and descriptions to maintain the database.
@@ Appendix F Prompts
F.1 Prompt for Decomposing User Requests
@@
Prompt for Decomposing User Requests Please help me break down a user request into multiple independent requirements, each including both positive and negative requirements. Return the results in the following format directly based on the **User Request** given, without writing any code. ### Output Format: Return a list, where each item is a dictionary representing an independent requirement, with the following key-value pairs: - **pos**: The positive requirement, representing what the user wants, excluding any negative requirements. - **neg**: The negative requirement, generally what the user does not want, dislikes, or refuses. All negative requirements must be captured in this field, for example, "non-spicy" should extract "spicy", "dont want crowded places" should extract "crowded", "hate noisy" should extract "noisy". - **mustsee**: Indicates whether this requirement represents a specific place name. If so, this field is true‘, otherwise, it is false‘. - **type**: Indicates whether the requirement is for a "place" or an "itinerary", with place having sub-types "place", "starting point", and "ending point". Overall, this field can have the values "place", "starting point", "ending point", or "itinerary". - Your return should be a list in the following format: [ {{ "pos": "positive requirement", (excluding negative requirements) "neg": "negative requirement" (whats not wanted, disliked, refused, not wanting to go or see, any negated requirement), "mustsee": true (whether its a must-see point, all specific places should be set to true), "type": "place" }}, ] - The **positive requirement** must not be empty, and it must not include any negative requirements. All negative requirements should be summarized in the value of the "neg" field. - Set to null in cases where there are no **negative requirements** for a specific place. - Sometimes users only describe what they do not want (negative requirements), in such cases, you should summarize a **positive requirement** based on the **negative requirement**. For example, if a user says dont want spicy food’, the output should include: "pos" corresponding to "food", "neg" corresponding to "spicy". - Independent requirements must have specific descriptions or demands to be considered a requirement, for example, "recommend a route" does not count as an independent requirement. - "mustsee" must be a specific place name, not a general term. - If a place is definitively a "starting point" or "ending point", then the value of the "type" field should be "starting point" or "ending point", respectively. "Starting points" and "ending points" are must-see points, with the "mustsee" field set to true. - A place can only be considered as a "starting point" or "ending point" if it is a specific attraction or location, and there can only be at most one "starting point" and one "ending point". - The return should not include any other content. ### Example Outputs: Example 1: User Request: "I want to start by visiting Sinan Mansions, then find something fun to do nearby, and I dont want it to be crowded" Output: [ {{ "pos": "Sinan Mansions", "neg": null, "mustsee": true, "type": "starting point" }}, {{ "pos": "fun places near Sinan Mansions", "neg": "crowded", "mustsee": false, "type": "place" }} ] ### mustsee Field Assignment Examples "mustsee" true for specific place names: "Hualian Supermarket", "Old Mac Coffee Shop", "Wukang Mansion", "Nanluoguxiang", "mustsee" false for general place names: "supermarket", "milk tea shop", "bar", "coffee", ### Output Guidelines - Return a list, each item in the list is a dictionary containing "pos", "neg", "mustsee", and "type" key-value pairs. - Return as a JSON List. - The list can be empty; if empty, just return a JSON list. - The output should not include any other information, ensuring it can be parsed by json.loads. ### User Request {user_req} ### Task Overview Your task is to analyze and break down the **User Request** into independent requirements and return them. 1. First, separate the different independent requirements, breaking down each into positive and negative requirements. 2. Positive requirements should only include what the user wants, and negative requirements should only include what the user does not want. 3. For each independent requirement, refer to **mustsee field assignment examples** to assign a value to the "mustsee" field, analyzing whether the **positive requirement** is a specific place name. If so, set "mustsee" to true, otherwise set it to false. 4. Refer to the **examples** and **output format** to complete the other fields. #### Notes: - Do not include duplicate independent requirements; ensure each independent requirement corresponds to different key points in the users needs. - "Itinerary" requirements should be for the whole itinerary, such as including several places, approximate time, etc., all others are place requirements. - The "type" field can only be one of ["place", "itinerary", "starting point", "ending point"]. - All attractions must be completely separated, such as "Nanluoguxiang and Drum Tower" must be split into "Nanluoguxiang" and "Drum Tower" as two requirements. Now, based on the **user Request**, refer to the **example outputs**, and return according to the **output guidelines** and **output format**.
F.2 Prompt for Indicating Travel Time of an Itinerary
@@
Prompt for Travel Time Indication Please play the role of a top AI Travel Time Planning Assistant. Your job is to determine the time needed for a days itinerary based on the user request. If the user request is empty, please default to ["4"]. ## Task Overview Your task is to return the required time for a days itinerary based on given the user request. If the user request mention specific time constraints for the route, return the itinerary time directly as per the users request, up to a maximum of 8 hours (return ["8"] if it exceeds 8 hours). Please return the itinerary time directly based on the user request, no need to write any code. ## User Request {user_reqs} ## Input-Output Examples - **Example 1**: - **User Request**: "Id like to visit a museum, enjoy authentic cuisine, and experience nightlife." - **Output**: ["8"] - **Example 2**: - **User Request**: "I want to tour historical buildings and take in the city views - **Output**: ["6"] - **Example 3** (In this example, the user specifies approximately **five hours** for the route): - **User Request**: "I plan to explore the Huangpu River and Yu Garden for about five hours." - **Output**: ["5"] ## Output Specifications - Return a list of length 1, containing a single integer representing the required itinerary duration (in hours). The value range is 1 to 8. - Return as a JSON List with only one element inside. - The list can be empty; please only return a JSON list. - Ensure your output contains no additional information and can be parsed by json.loads. Now, based on the **User Request**, return the time required for a days itinerary according to the **Output Specifications**.
@@
F.3 Prompt for Identifying the Start POI